OpenStudy (anonymous):

In October, Greg and Thomas had the same amount of money in their savings accounts. In November, Greg deposited $120 into his account. Thomas increased the money in his account by 20%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in October?

OpenStudy (texaschic101):

x = starting balance x + 120 = x + 0.20x x + 120 = 1.20x 120 = 1.20x - x 120 = .20x 120/.20 = x 600 = x check.. x + 120 = x + 0.20x 600 + 120 = 600 + 0.20(600) 720 = 720 (correct) In October, they both had $ 600. In November, they both ended up with $ 720

OpenStudy (anonymous):

thank youu